# Using the Nucleolus for Incentive Allocation in Vertical Federated Learning

### v2

@@ Abstract
Vertical federated learning (VFL) is a promising approach for collaboratively training machine learning models using private data partitioned vertically across different parties. Ideally in a VFL setting, the active party (party possessing features of samples with labels) benefits by improving its machine learning model through collaboration with some passive parties (parties possessing additional features of the same samples without labels) in a privacy preserving manner. However, motivating passive parties to participate in VFL can be challenging. In this paper, we focus on the problem of allocating incentives to the passive parties by the active party based on their contributions to the VFL process. We address this by formulating the incentive allocation problem as a bankruptcy game, a concept from cooperative game theory. Using the Talmudic division rule, which leads to the Nucleolus as its solution, we ensure a fair distribution of incentives. We evaluate our proposed method on synthetic and real-world datasets and show that it ensures fairness and stability in incentive allocation among passive parties who contribute their data to the federated model. Additionally, we compare our method to the existing solution of calculating Shapley values and show that our approach provides a more efficient solution with fewer computations.
@@ 1 Introduction
Motivating data owners to participate in federated learning is essential for its success. Incentives encourage data owners to contribute data, improving model performance with diverse and representative data sources. Shi et al. [13] proposed a framework in an HFL setting, using Shapley values to allocate incentives based on contributions, with a reputation-driven reward allocation policy. Similarly, Zhang et al. [24] proposed a reward system in HFL where clients bid prices and a selector chooses participants based on contribution quality, with reputation and payment dependent on the quality of contributions. Incentive mechanisms differ between HFL and VFL due to the different roles of data owners and model goals. Unlike an HFL setting, there are two types of data owners or parties involved in the federation in VFL: : (1) Passive/Host Party, owning data (features) without labels for model training; and (2) Active/Guest Party, holding labeled data for training models using host data. Hence, the active party in VFL substantially benefits from additional data provided by the passive parties, necessitating a motivating mechanism for the latter to contribute data. However, most of the existing approaches for incentive allocation in federated learning deal with horizontally partitioned data, where the data is partitioned based on the samples. In contrast, vertically partitioned data is partitioned based on the features or attributes. To the best of our knowledge, no existing studies have formulated the incentive allocation problem in VFL as a bankruptcy game, which offers a unique perspective for addressing this challenge in certain scenarios. Our main contributions in this paper include:
@@ 3 Method
In this section, we propose our approach to incentive allocation in VFL formulating the problem as a bankruptcy game. The choice of a bankruptcy game is predicated on its suitability for scenarios where resources—in this case, incentives—must be distributed among parties who have claims (i.e., contributions to the learning process) that exceed the available resources. This framework is particularly applicable for VFL settings, where the active party, benefiting directly from the collaborative learning process, must determine how to fairly distribute a finite pool of incentives among multiple passive parties, each contributing additional data features that improve the learning model. For completeness we present here the definitions of the bankruptcy game.
@@ 3.1 Reminder On Bankruptcy Game
A bankruptcy game, based on the bankruptcy problem, arises when an entity’s assets (denoted by $E$) are insufficient to meet all creditors’ claims, $c_{i}$, for $i=1,...,n$. The game is defined by a set of players (claimants) $C=\{1,2,\ldots,n\}$, each with a claim $c_{i}$ on the estate, where typically $\sum_{i=1}^{n}c_{i}>E$. In this context, the coalition value for a subset of claimants $S\subseteq C$ is determined by computing the remaining estate after fulfilling the claims of claimants not involved in the that coalition.

### v3

@@ Abstract
Vertical federated learning (VFL) is a promising approach for collaboratively training machine learning models using private data partitioned vertically across different parties. Ideally in a VFL setting, the active party (party possessing features of samples with labels) benefits by improving its machine learning model through collaboration with some passive parties (parties possessing additional features of the same samples without labels) in a privacy-preserving manner. However, motivating passive parties to participate in VFL can be challenging. In this paper, we focus on the problem of allocating incentives to the passive parties by the active party based on their contributions to the VFL process. We address this by formulating the incentive allocation problem as a bankruptcy game, a concept from cooperative game theory. Using the Talmudic division rule, which leads to the Nucleolus as its solution, we ensure a fair distribution of incentives. We evaluate our proposed method on synthetic and real-world datasets and show that it ensures fairness and stability in incentive allocation among passive parties who contribute their data to the federated model. Additionally, we compare our method to the existing solution of calculating Shapley values and show that our approach provides a more efficient solution with fewer computations.
@@ 1 Introduction
Motivating data owners to participate in federated learning is essential for its success. Incentives encourage data owners to contribute data, improving model performance with diverse and representative data sources. Shi et al. [13] proposed a framework in an HFL setting, using Shapley values to allocate incentives based on contributions, with a reputation-driven reward allocation policy. Similarly, Zhang et al. [24] proposed a reward system in HFL where clients bid prices and a selector chooses participants based on contribution quality, with reputation and payment dependent on the quality of contributions. Incentive mechanisms differ between HFL and VFL due to the different roles of data owners and model goals. Unlike an HFL setting, there are two types of data owners or parties involved in the federation in VFL: (1) Passive/Host Party, owning data (features) without labels for model training; and (2) Active/Guest Party, holding labeled data for training models using host data. Hence, the active party in VFL substantially benefits from additional data provided by the passive parties, necessitating a motivating mechanism for the latter to contribute data. However, most of the existing approaches for incentive allocation in federated learning deal with horizontally partitioned data, where the data is partitioned based on the samples. In contrast, vertically partitioned data is partitioned based on the features or attributes. To the best of our knowledge, no existing studies have formulated the incentive allocation problem in VFL as a bankruptcy game, which offers a unique perspective for addressing this challenge in certain scenarios. Our main contributions in this paper include:
@@ 3 Method
In this section, we propose our approach to incentive allocation in VFL formulating the problem as a bankruptcy game. The choice of a bankruptcy game is predicated on its suitability for scenarios where resources—in this case, incentives—must be distributed among parties who have claims (i.e., contributions to the learning process) that exceed the available resources. This framework is particularly applicable for VFL settings, where the active party, benefiting directly from the collaborative learning process, must determine how to fairly distribute a finite pool of incentives among multiple passive parties, each contributing additional data features that improve the learning model. For completeness, we present here the definitions of the bankruptcy game.
@@ 3.1 Reminder On Bankruptcy Game
A bankruptcy game, based on the bankruptcy problem, arises when an entity’s assets (denoted by $E$) are insufficient to meet all creditors’ claims, $c_{i}$, for $i=1,...,n$. The game is defined by a set of players (claimants) $C=\{1,2,\ldots,n\}$, each with a claim $c_{i}$ on the estate, where typically $\sum_{i=1}^{n}c_{i}>E$. In this context, the coalition value for a subset of claimants $S\subseteq C$ is determined by computing the remaining estate after fulfilling the claims of claimants not involved in that coalition.